Find the equivalent resistance of the circuit. 400 20 a 80 w 30 W O a 85 1700 Ob. c. 80.50 d. 1000
Answer #1

For resistors connected in series,

Equivalent resistance, R_{eq}=R_{1}+R_{2}+R_{3}+............

and for resistors connected in parallel,

equivalent resistance, \frac{1}{R_{eq}}=\frac{1}{R_{1}}+\frac{1}{R_{2}}+\frac{1}{R_{3}}+.........

===============

Consider the network

Resistors 40Ω αnd 20Ω are connected in series

R_{s}=40\Omega +20\Omega

{\color{Red} R_{s}=60\Omega }

----------

Resistors R_{s}\ and\ 30\Omega are connected in parallel

\frac{1}{R_{p}}=\frac{1}{R_{s}}+\frac{1}{30\Omega }

\frac{1}{R_{p}}=\frac{1}{60\Omega }+\frac{1}{30\Omega }

\frac{1}{R_{p}}=\frac{1}{60}+\frac{1}{30}=\frac{1}{60}+\frac{2}{60}=\frac{1+2}{60}=\frac{3}{60}=\frac{1}{20}

\frac{1}{R_{p}}=\frac{1}{20}

{\color{Red} R_{p}=20\Omega}

===============

Resistors 80\Omega \ and\ R_{p} are connected in series

R_{eq}=80\Omega +R_{P}

R_{eq}=80\Omega +20\Omega

{\color{Red} R_{eq}=100\Omega }

ANSWER: Option d. 100\ \Omega

===========================